OpenStudy (anonymous):

what is the equivlaent expression to csc (-pi/3)

OpenStudy (jonnyvonny):

Csc(x) is the same as 1/sin(x). If that's not what they want, 1/(sin(-pi/3))=\[1/(\sqrt{3}/2)\rightarrow 2/\sqrt{3}\] Which is not part of the unit circle, so I believe it was the first statement, lol.